what is a domain? and how would i figure this out? is there an easy way to solve this?? can you please help thank you very much

Determine the domain of each relation.
28. y = \/ x-2
30. y = 4-x^2

Answer by venugopalramana(3286)

OK....DOMAIN I SHALL TELL YOU..
THE VALUES X CAN TAKE TO MAKE THE FUNCTION REAL IS
DOMAIN.
EX.1
F(X)=X^2+1....X CAN TAKE ANY REAL VALUE..SO DOMAIN IS
ALL REAL NUMBERS.
IN SET NOTATION WE WRITE IT AS (-INFINITY,+INFINITY)
EX.2
G(X)=SQRT(X)
NOW IF X IS NEGATIVE THE FUNCTION BECOMES
IMAGINARY..HENCE X CANNOT BE NEGATIVE.
SO WE SAY DOMAIN OF X IS ALLREAL VALUES GREATER THAN
OR EQUAL TO ZERO.
IN SET NOTATION WE WRITE IT AS (0,INFINITY)
YOUR PROBLEM
Y = SQRT(X-2).....X-2>=0
X>=2
DOMAIN IS +2 TO INFINITY...[2,INFINITY)...[ INDICATES IT CAN EQUAL 2